# Distillation Matters: Empowering Sequential Recommenders to Match the Performance of Large Language Models

### v1

@@ Abstract.
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/18/06ccs: Information systems Recommender systems
@@ 1. INTRODUCTION
Large Language Models (LLMs) have showcased remarkable capabilities in content comprehension, generation, and semantic reasoning (Brown et al., 2020; Achiam et al., 2023; Touvron et al., 2023). Recently, LLMs have sparked a surge of interest within the domain of Recommender Systems (RS). Various research efforts have been devoted to harnessing LLMs to augment traditional recommendation models, serving as encoders for user/item features or as supplementary knowledge bases (Hou et al., 2022, 2023; Ren et al., 2023; Rajput et al., 2024). To fully exploit the reasoning ability of LLMs in RS, another line of research is to directly prompt or fine-tune LLMs to function as specialized recommenders. Owing to their inherent semantic reasoning capabilities, these LLM-based recommendation methods have achieved impressive performance. For instance, as shown in Table  1, the representative model BIGRec (Bao et al., 2023a) has demonstrated approximately 41.44% improvements on average over the state-of-the-art conventional sequential model (i.e., DROS (Yang et al., 2023)) on the typical Amazon Games and Toys datasets.
@@
To tackle this challenge, we propose employing knowledge distillation (KD) for acceleration — i.e., distilling the knowledge from a complex LLM-based recommendation model (teacher) to a lightweight conventional recommendation model (student). KD has been successfully applied in many domains (Hinton et al., 2015; Adriana et al., 2015; Chen et al., 2023), and has been proven effective in transferring knowledge from a large model to a smaller one. This strategy could capitalize on the effectiveness of LLM-based recommenders while maintaining low inference costs. It also potentially integrates the capabilities of conventional models in capturing collaborative signals with the semantic reasoning prowess of LLMs. While the ideal is promising, distillation is non-trivial due to the fundamentally different mechanisms between the teacher and student models. LLMs primarily rely on content understanding and capturing semantic correlations for making recommendations, whereas conventional models depend on collaborative signals derived from users’ historical behaviors. This divergence introduces several challenges:
@@ 2.2. Brief on BIGRec
Recently LLM-based recommendation attracts great attention. Predominantly, this body of work formulates the recommendation task using natural language prompts and employs large language models to generate personalized recommendations (Hou et al., 2024). This study simply take the representative model BIGRec (Bao et al., 2023a) for empirical analysis. The selection of BIGRec is justified not only by its availability as an open-source tool but also by its demonstrated effectiveness. Furthermore, BIGRec embodies the fundamental elements of LLM-based recommendation and many methods can be considered as further extensions of such paradigm (Sun et al., 2024; Wang et al., 2024d). It is also noteworthy that BIGRec has been employed by recent studies (Lin et al., 2024, 2023b) as a representative model for analysis.
To be specific, BIGRec organizes users’ historical behaviors in natural language and employs instruction-tuning to fine-tune LLMs, as illustrated in Figure 1. During the inference stage, BIGRec generates item descriptions (e.g., titles) for recommendations. Considering that these descriptions may not always correspond to existing items, BIGRec incorporates a grounding paradigm that matches generated item descriptions to existing items based on content similarity. Formally, let $\mathbf{z}_{d}$ and $\mathbf{z}_{i}$ represent the token embeddings of generated descriptions and the descriptions of item $i$, respectively. BIGRec computes their L2 distance for grounding as follows:
(4) $d_{\mathbf{s}i}=||\mathbf{z}_{d_{\mathbf{s}}}-\mathbf{z}_{i}||^{2}$
@@ 3.1. Motivations
Table 2. The ratio of cases where BIGRec outperforms DROS to cases where BIGRec underperforms DROS.
@@ 3.2. Proposed Distillation Strategy: DLLM2Rec
where $\mathcal{O}^{T}$ denotes the Top-K recommendations returned by the teacher model. In this work, we choose $K=10$ as a default value, but this can be tuned for optimal performance. The objective is straightforward — we select the highly ranked items from the teacher as a positive to guide the learning of the student, so that these candidate items can also be recommended by the student. However, given that the teacher’s recommendations may not always be beneficial, we introduce an importance weight that considers the following three aspects:
@@
$d_{\mathbf{s}i^{*}}=||\mathbf{z}_{d_{\mathbf{s}}}-\mathbf{z}_{i^{*}}||^{2}$
where $d_{\mathbf{s}i^{*}}$ measures the embedding distance between the generated item description $\mathbf{z}_{d_{\mathbf{s}}}$ and the target ground-truth item $\mathbf{z}_{i^{*}}$, where the embeddings can be generated via LLMs encoder. A smaller distance suggests a higher quality of the generated description as it aligns more closely with the targets. Conversely, a larger gap suggests lower confidence, indicating that LLMs may risk generating incorrect or nonsensical information.
@@
(9) $\frac{\partial\mathcal{L}_{d}}{\partial\hat{y}_{\mathbf{s}i}}=w_{\mathbf{s}i}\sigma(1-\hat{y}_{\mathbf{s}i})$
@@ 4.1.1. Datasets
Three conventional datasets: Amazon Video Games, MovieLens-10M, and Amazon Toys and Games were utilized in our experiments. These datasets include user behavior sequences and item content. For fair comparisons, we closely adhered to the preprocessing methods used in recent work (Yang et al., 2023; Bao et al., 2023a). We organized the interaction sequences in ascending order of timestamps to partition each dataset into training, validation, and testing sets with ratios of 8:1:1. Given that MovieLens-10M contains an excessive number of sequences, which could not be processed by LLM-based recommenders, we sampled 100,000 sequences for training and 10,000 for testing, the sampling strategy also adopted by (Bao et al., 2023a). The dataset statistics are presented in Table 4.
@@ 4.1.2. Baselines
Knowledge Distillation Strategies for Recommendation: Hint (Adriana et al., 2015), HTD (Kang et al., 2021) are two representative methods that distill knowledge through teacher embeddings; RD (Tang and Wang, 2018b), CD (Lee et al., 2019), RRD (Kang et al., 2020), DCD (Lee and Kim, 2021) and UnKD (Chen et al., 2023) distill information from the teachers’ ranking lists. Readers may refer to the related work for more details about these strategies.
@@ 4.1.3. Evaluation Metrics
We employed two widely-used metrics HR@K and NDCG@K to evaluate performance. Here we simply sest K to 20 as recent work (He et al., 2020), and observed similar results with other choices of $K$.
@@ 4.1.4. Implementation Details
All methods are implemented with PyTorch and run on 4 Nvidia A800 GPUs. We set $\beta=1.0,\gamma_{p}=0.3,\gamma_{c}=0.5,\gamma_{o}=0.1$ across all datasets, as these settings were found sufficient to generate good performance, although fine-tuning could further enhance model performance. The influence of these hyperparameters on model performance is also presented in Figure 2. Adam (Kingma and Ba, 2014) was used as the optimizer with a tuned learning rate of 0.001, a batch size of 256, and weight decay tuned in {1e-4, 1e-5, 1e-6, 1e-7, 0}, $\lambda_{d}$ in {0.1, 0.2, …, 1.0}. We set the embedding size to 64 and the dropout ratio to 0.1.
@@ 5.1. LLMs for Recommendation System
LLMs as recommenders. Initial efforts explored the zero-shot capabilities of LLMs in recommendation by structuring the recommendation tasks as language prompts (Gao et al., 2023; Hou et al., 2024; Wang and Lim, 2023; Liu et al., 2023; Wang et al., 2024b). Subsequently, to adapt LLMs to recommendation tasks, instruction-tuning or fine-tuning has been widely adopted, showing promising results (Bao et al., 2023a, b; Li et al., 2023a; Zhang et al., 2023; Li et al., 2024b; Wang et al., 2024a; Hu et al., 2024; Shi et al., 2024; Li et al., 2023b; Zhang et al., 2024b; Li et al., 2024a). This research primarily focuses on how to enhance LLMs to better suit recommendation tasks. For instance, some studies aim to minimize the semantic gap between recommendations and natural language (Bao et al., 2023a; Zhu et al., 2023; Hou et al., 2023; Zheng et al., 2023; Wang et al., 2024c; Liao et al., 2023; Xu et al., 2024; Cao et al., 2024; Jiang et al., 2024; Tan et al., 2024), such as the approach taken by BIGRec (Bao et al., 2023a), which uses a grounding strategy to map LLM descriptions to recommended items. Others focus on improving LLMs’ ability to model long-sequence interactions (Geng et al., 2024; Zheng et al., 2024; Lin et al., 2023a), and some attempt to reduce training overhead (Lin et al., 2024). While effective, these methods often suffer from significant inference inefficiency, limiting their practical application. Although some studies have tried to mitigate inference latency through pre-storage techniques (Geng et al., 2024) or knowledge distillation (Wang et al., 2024d), the gain of (Geng et al., 2024) is generally modest and (Wang et al., 2024d) is utilized to distill a huge LLMs (e.g., GPT-115B) to a relatively smaller LLMs (e.g., Llama-7B). Even small Llama-7B is hard to deploy in practical.
@@
LLMs enhancing conventional recommenders. Existing methods mainly employ LLMs as supplementary knowledge bases (Yang et al., 2024; Xi et al., 2023; Qin et al., 2024) or as encoders for users/items (Ren et al., 2023; Wei et al., 2024; Wang et al., 2023) to improve conventional recommenders. For instance, KAR (Xi et al., 2023) utilizes LLMs as external knowledge bases to better profile users and items within the recommender system. RLMRec (Ren et al., 2023) encodes user and item profiles into semantic representations and aligns embeddings generated by conventional models with those from LLMs; Wei et al. (Wei et al., 2024) exploit LLMs to uncover new relationships within graphs. However, compared to direct LLM-based recommenders, these methods do not fully leverage the semantic reasoning capabilities of LLMs for making recommendations. CSRec (Yang et al., 2024) incorporates the common sense extrcted from LLM to alleviate the data sparsity issue.
To capitalize on the reasoning ability of teachers, some efforts have attempted to use chains of thought (CoT) data generated by LLMs to enhance conventional models (Sun et al., 2024; Wang et al., 2024d). However, given the relatively simple architecture of conventional models, it is notably difficult for these systems to assimilate the complex knowledge from CoT data. Furthermore, CoT is typically used as a feature for a sequence, indicating that CoT for a new sequence should be produced during inference, which would be time-consuming. While some approaches attempt to retrieve similar CoT from other user-item pairs (Sun et al., 2024), such approximations may hurt accuracy.
@@ 5.2. Sequential Recommendation
Sequential recommendation (Sun et al., 2019; Xie et al., 2022; Chen et al., 2018; Chang et al., 2021; Li et al., 2020) takes into account the sequence or order of interactions to predict what a user might prefer next. Existing Sequential RS use sequence generation models, such as RNNs or Transformers, to model user interaction sequences. For example, GRU4Rec (Hidasi et al., 2015) employs the GRU to handle session-based data, while Caser (Tang and Wang, 2018a) uses CNN to model interaction data on multiple levels. SASRec (Kang and McAuley, 2018) introduces attention mechanism to automatically learn the weights of different interaction items, and DROS leverage distributional robust optimization in sequential recommendation and achieves state-of-the-art performance. Most current LLM for RS methods also adopt the setting of sequential recommendation (Bao et al., 2023a, b). The readers may refer to the excellent survey (Fang et al., 2020; Wang et al., 2019) for more details.
@@ 5.3. Knowledge Distillation in RS
Knowledge distillation (KD) is a promising model compression technique that transfers the knowledge from a large teacher model into the target compact student model (Gou et al., 2021), and they have been widely applied in recommendation systems to reduce inference latency. RD (Tang and Wang, 2018b) treated the top-N ranked items as positive for training a student model; CD (Lee et al., 2019) utilized soft labels to create positive and negative distillation instances; Soft labels also considered by RRD (Kang et al., 2020) to create the list-wise distillation loss function; DCD (Lee and Kim, 2021) built distillation loss on both user-side and item-side; UnKD (Chen et al., 2023) addresses popularity bias in distillation. The hidden knowledge among the middle layer of teachers are also considered in some methods. For example, Hint (Hinton et al., 2015) and RRD (Kang et al., 2020) extracted knowledge of teachers’ embedding via Fitnet and expert network; HTD (Kang et al., 2021) distilled the topological knowledge with the relations in the teacher embedding space. Some researcher also study to distill knowledge from a huge LLMs (e.g., ChatGPT) to a relatively smaller LLMs (e.g., LLaMA-7B) in the recommendation scenarios. Besides model compression, KD is also used to integrate knowledge among different models (Tao et al., 2022; Xu et al., 2022; Zhu et al., 2020). For example, some researchers (Zhu et al., 2020) consider to integrate knowledge from multiply pre-trained models into the student. To the best of our knowledge, the study of distilling LLM-based recommenders into conventional recommenders remains untouched.

### v3

@@ Abstract.
journalyear: 2024copyright: acmlicensedconference: 18th ACM Conference on Recommender Systems; October 14–18, 2024; Bari, Italybooktitle: 18th ACM Conference on Recommender Systems (RecSys ’24), October 14–18, 2024, Bari, Italydoi: 10.1145/3640457.3688118isbn: 979-8-4007-0505-2/24/10submissionid: 120ccs: Information systems Recommender systems
@@ 1. INTRODUCTION
Large Language Models (LLMs) have showcased remarkable capabilities in content comprehension, generation, and semantic reasoning (Brown et al., 2020; Achiam et al., 2023; Touvron et al., 2023). Recently, LLMs have sparked a surge of interest within the domain of Recommender Systems (RS). Various research efforts have been devoted to harnessing LLMs to augment traditional recommendation models, serving as encoders for user/item features or as supplementary knowledge bases (Hou et al., 2022, 2023; Ren et al., 2024; Rajput et al., 2024). To fully exploit the reasoning ability of LLMs in RS, another line of research is to directly prompt or fine-tune LLMs to function as specialized recommenders. Owing to their inherent semantic reasoning capabilities, these LLM-based recommendation methods have achieved impressive performance. For instance, as shown in Table  1, the representative model BIGRec (Bao et al., 2023a) has demonstrated approximately 41.44% improvements on average over the state-of-the-art conventional sequential model (i.e., DROS (Yang et al., 2023)) on the typical Amazon Games and Toys datasets.
@@
To tackle this challenge, we propose employing knowledge distillation (KD) for acceleration — i.e., distilling the knowledge from a complex LLM-based recommendation model (teacher) to a lightweight conventional recommendation model (student). KD has been successfully applied in many domains (Hinton et al., 2015; Adriana et al., 2015; Chen et al., 2023a), and has been proven effective in transferring knowledge from a large model to a smaller one. This strategy could capitalize on the effectiveness of LLM-based recommenders while maintaining low inference costs. It also potentially integrates the capabilities of conventional models in capturing collaborative signals with the semantic reasoning prowess of LLMs. While the ideal is promising, distillation is non-trivial due to the fundamentally different mechanisms between the teacher and student models. LLMs primarily rely on content understanding and capturing semantic correlations for making recommendations, whereas conventional models depend on collaborative signals derived from users’ historical behaviors. This divergence introduces several challenges:
@@ 2.2. Brief on BIGRec
Recently LLM-based recommendation attracts great attention. Predominantly, this body of work formulates the recommendation task using natural language prompts and employs large language models to generate personalized recommendations (Hou et al., 2024). This study simply take the representative model BIGRec (Bao et al., 2023a) for empirical analysis. The selection of BIGRec is justified not only by its availability as an open-source tool but also by its demonstrated effectiveness. Furthermore, BIGRec embodies the fundamental elements of LLM-based recommendation and many methods can be considered as further extensions of such paradigm (Sun et al., 2024; Wang et al., 2024f). It is also noteworthy that BIGRec has been employed by recent studies (Lin et al., 2024c, 2023) as a representative model for analysis.
To be specific, BIGRec organizes users’ historical behaviors in natural language and employs instruction-tuning to fine-tune LLMs, as illustrated in Figure 1. During the inference stage, BIGRec generates item descriptions (e.g., titles) for recommendations. Considering that these descriptions may not always correspond to existing items, BIGRec incorporates a grounding paradigm that matches generated item descriptions to existing items based on content similarity. Formally, let $\mathbf{z}_{g_{\mathbf{s}}}$ and $\mathbf{z}_{i}$ represent the token embeddings of generated descriptions and the descriptions of item $i$, respectively. BIGRec computes their L2 distance for grounding as follows:
(4) $d_{\mathbf{s}i}=||\mathbf{z}_{g_{\mathbf{s}}}-\mathbf{z}_{i}||^{2}$
@@ 3.1. Motivations
Table 2. The ratio of cases where BIGRec outperforms DROS to cases where BIGRec underperforms DROS on NDCG@20.
@@ 3.2. Proposed Distillation Strategy: DLLM2Rec
where $\mathcal{O}^{T}$ denotes the Top-K recommendations returned by the teacher model, and $w_{si}$ denotes the distillation weight. In this work, we choose $K=10$ as a default value, but this can be tuned for optimal performance. The objective is straightforward — we select the highly ranked items from the teacher as a positive to guide the learning of the student, so that these candidate items can also be recommended by the student. However, given that the teacher’s recommendations may not always be beneficial, we introduce an importance weight that considers the following three aspects:
@@
$d_{\mathbf{s}i^{*}}=||\mathbf{z}_{{g_{\mathbf{s}}}}-\mathbf{z}_{i^{*}}||^{2}$
where $d_{\mathbf{s}i^{*}}$ measures the embedding distance between the generated item description $\mathbf{z}_{g_{\mathbf{s}}}$ and the target ground-truth item $\mathbf{z}_{i^{*}}$, where the embeddings can be generated via LLMs encoder. A smaller distance suggests a higher quality of the generated description as it aligns more closely with the targets. Conversely, a larger gap suggests lower confidence, indicating that LLMs may risk generating incorrect or nonsensical information.
@@
(9) $\frac{\partial\mathcal{L}_{d}}{\partial\hat{y}_{\mathbf{s}i}}=w_{\mathbf{s}i}\sigma(-\hat{y}_{\mathbf{s}i})$
@@ 4.1.1. Datasets
Three conventional datasets: Amazon Video Games, MovieLens-10M, and Amazon Toys and Games were utilized in our experiments  [^1] https://jmcauley.ucsd.edu/data/amazon/, [^2] https://grouplens.org/datasets/movielens/10m/. These datasets include user behavior sequences and item content. For fair comparisons, we closely adhered to the preprocessing methods used in recent work (Yang et al., 2023; Bao et al., 2023a). We organized the interaction sequences in ascending order of timestamps to partition each dataset into training, validation, and testing sets with ratios of 8:1:1. Given that MovieLens-10M contains an excessive number of sequences, which could not be processed by LLM-based recommenders, we sampled 100,000 sequences for training and 10,000 for testing, the sampling strategy also adopted by (Bao et al., 2023a). The dataset statistics are presented in Table 4.
@@ 4.1.2. Baselines
Knowledge Distillation Strategies for Recommendation: Hint (Adriana et al., 2015), HTD (Kang et al., 2021) are two representative methods that distill knowledge through teacher embeddings; RD (Tang and Wang, 2018b), CD (Lee et al., 2019), RRD (Kang et al., 2020), DCD (Lee and Kim, 2021) and UnKD (Chen et al., 2023a) distill information from the teachers’ ranking lists. Readers may refer to the related work for more details about these strategies.
@@ 4.1.3. Evaluation Metrics
We employed two widely-used metrics HR@K and NDCG@K to evaluate performance. Here we simply set K to 20 as recent work (He et al., 2020), and observed similar results with other choices of $K$.
@@ 4.1.4. Implementation Details
All methods are implemented with PyTorch and run on 4 Nvidia A800 GPUs. We set $\beta=1.0,\gamma_{p}=0.3,\gamma_{c}=0.5,\gamma_{o}=0.1$ across all datasets, as these settings were found sufficient to generate good performance, although fine-tuning could further enhance model performance. The influence of these hyperparameters on model performance is also presented in Figure 2. Adam (Kingma and Ba, 2014) was used as the optimizer with a tuned learning rate of 0.001, a batch size of 256, and weight decay tuned in {1e-4, 1e-5, 1e-6, 1e-7, 0}, $\lambda_{d}$ in {0.1, 0.2, …, 1.0}. We set the embedding size to 64 and the dropout ratio to 0.1. Our code is available on  [^3] https://github.com/istarryn/DLLM2Rec.
@@ 4.2. Performance Comparison (RQ1)
Table 7. Overlapping ratio on Top-20 items.
Datasets Before-distillation Post-distillation
Games 3.99% 10.88%
MovieLens 4.06% 10.15%
Toys 1.05% 14.56%
Besides, we conducted additional experiments to determine the ratio of overlapping items between the teacher and student models. As shown in Table 7, the post-distillation student model can effectively assimilate the teacher knowledge. Furthermore, it is important to note that the post-distillation student model may not entirely replicate the teacher’s recommendations, given that the potential unreliability of teacher knowledge and the inherent teacher-student capacity gap.
@@ 5.1. LLMs for Recommendation System
LLMs as recommenders. Initial efforts explored the zero-shot capabilities of LLMs in recommendation by structuring the recommendation tasks as language prompts (Gao et al., 2023; Hou et al., 2024; Wang and Lim, 2023; Liu et al., 2023; Wang et al., 2024d). Subsequently, to adapt LLMs to recommendation tasks, instruction-tuning or fine-tuning has been widely adopted, showing promising results (Bao et al., 2023a, b; Li et al., 2023a; Zhang et al., 2023; Li et al., 2024b; Wang et al., 2024b; Hu et al., 2024; Shi et al., 2024; Li et al., 2023b; Zhang et al., 2024b; Li et al., 2024a). This research primarily focuses on how to enhance LLMs to better suit recommendation tasks. For instance, some studies aim to minimize the semantic gap between recommendations and natural language (Bao et al., 2023a; Zhu et al., 2024; Hou et al., 2023; Zheng et al., 2023; Wang et al., 2024e; Liao et al., 2024; Xu et al., 2024; Cao et al., 2024; Jiang et al., 2024; Tan et al., 2024), such as the approach taken by BIGRec (Bao et al., 2023a), which uses a grounding strategy to map LLM descriptions to recommended items. Others focus on improving LLMs’ ability to model long-sequence interactions (Geng et al., 2024; Zheng et al., 2024; Lin et al., 2024b), identify noisy items (Wang et al., 2024c) and some attempt to reduce training overhead (Lin et al., 2024c). While effective, these methods often suffer from significant inference inefficiency, limiting their practical application. Although some studies have tried to mitigate inference latency through pre-storage techniques (Geng et al., 2024) or knowledge distillation (Wang et al., 2024f), the gain of (Geng et al., 2024) is generally modest and (Wang et al., 2024f) is utilized to distill a huge LLMs (e.g., GPT-115B) to a relatively smaller LLMs (e.g., Llama-7B). Even small Llama-7B is hard to deploy in practical.
@@
LLMs enhancing conventional recommenders. Existing methods mainly employ LLMs as supplementary knowledge bases (Yang et al., 2024; Xi et al., 2023; Qin et al., 2024) or as encoders for users/items (Ren et al., 2024; Wei et al., 2024; Wang et al., 2023) to improve conventional recommenders. For instance, KAR (Xi et al., 2023) utilizes LLMs as external knowledge bases to better profile users and items within the recommender system. RLMRec (Ren et al., 2024) encodes user and item profiles into semantic representations and aligns embeddings generated by conventional models with those from LLMs; Wei et al. (Wei et al., 2024) exploit LLMs to uncover new relationships within graphs. However, compared to direct LLM-based recommenders, these methods do not fully leverage the semantic reasoning capabilities of LLMs for making recommendations. CSRec (Yang et al., 2024) incorporates the common sense extracted from LLM to alleviate the data sparsity issue.
To capitalize on the reasoning ability of teachers, some efforts have attempted to use chains of thought (CoT) data generated by LLMs to enhance conventional models (Sun et al., 2024; Wang et al., 2024f). However, given the relatively simple architecture of conventional models, it is notably difficult for these systems to assimilate the complex knowledge from CoT data. Furthermore, CoT is typically used as a feature for a sequence, indicating that CoT for a new sequence should be produced during inference, which would be time-consuming. While some approaches attempt to retrieve similar CoT from other user-item pairs (Sun et al., 2024), such approximations may hurt accuracy.
@@ 5.2. Sequential Recommendation
Sequential recommendation (Sun et al., 2019; Xie et al., 2022; Chen et al., 2018; Chang et al., 2021; Li et al., 2020) takes into account the sequence or order of interactions to predict what a user might prefer next. Existing Sequential RS use sequence generation models, such as RNNs (Hidasi et al., 2015) or Transformers (Kang and McAuley, 2018; Chen et al., 2024; Yang et al., 2023), to model user interaction sequences. For example, GRU4Rec (Hidasi et al., 2015) employs the GRU to handle session-based data, while Caser (Tang and Wang, 2018a) uses CNN to model interaction data on multiple levels. SASRec (Kang and McAuley, 2018) introduces attention mechanism to automatically learn the weights of different interaction items, and DROS (Yang et al., 2023) leverage distributional robust optimization in sequential recommendation and achieves state-of-the-art performance. Some other methods focus on the intrinsic biases (Lin et al., 2024a; Chen et al., 2023b, 2021; Zhao et al., 2022) and distribution shifts (Wang et al., 2024a) within RS. Most current LLM for RS methods also adopt the setting of sequential recommendation (Bao et al., 2023a, b). The readers may refer to the excellent survey (Fang et al., 2020; Wang et al., 2019) for more details.
@@ 5.3. Knowledge Distillation in RS
Knowledge distillation (KD) is a promising model compression technique that transfers the knowledge from a large teacher model into the target compact student model (Gou et al., 2021), and they have been widely applied in recommendation systems to reduce inference latency. RD (Tang and Wang, 2018b) treated the top-N ranked items as positive for training a student model; CD (Lee et al., 2019) utilized soft labels to create positive and negative distillation instances; Soft labels also considered by RRD (Kang et al., 2020) to create the list-wise distillation loss function; DCD (Lee and Kim, 2021) built distillation loss on both user-side and item-side; UnKD (Chen et al., 2023a) addresses popularity bias in distillation. The hidden knowledge among the middle layer of teachers are also considered in some methods. For example, Hint (Hinton et al., 2015) and RRD (Kang et al., 2020) extracted knowledge of teachers’ embedding via Fitnet and expert network; HTD (Kang et al., 2021) distilled the topological knowledge with the relations in the teacher embedding space. Some researcher also study to distill knowledge from a huge LLMs (e.g., ChatGPT) to a relatively smaller LLMs (e.g., LLaMA-7B) in the recommendation scenarios. Besides model compression, KD is also used to integrate knowledge among different models (Tao et al., 2022; Xu et al., 2022; Zhu et al., 2020). For example, some researchers (Zhu et al., 2020) consider to integrate knowledge from multiply pre-trained models into the student. To the best of our knowledge, the study of distilling LLM-based recommenders into conventional recommenders remains untouched.